OpenStudy (anonymous):

If x is not equal to 0 and x is inversely proportional to y, which of the following is directly proportional to 1/x sq.?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

If x is inversely proportional to y, then y is inversely proportional to x. You can write this as either y=k/x or x=k/y. Maybe substitute k/y for x and square that.

OpenStudy (anonymous):

Not sure if I am correct, but let x = k/y. k is a constant so you don't have to worry about it. Now reciprocal both side then square it.

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

1/x=y/k?

OpenStudy (anonymous):

yeah, now square both sides.

OpenStudy (anonymous):

Ignore the constant, we want to find what function of y is directly proportional to 1/x².

OpenStudy (anonymous):

1/x squared=y squared/k squared?

OpenStudy (anonymous):

Yeah, ignoring the constant, y² is directly proportional to 1/x², I believe.

OpenStudy (anonymous):

By the way, is your question the multiple choice question?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

1/x sq.=y sq./ k sq.

OpenStudy (anonymous):

k sq./x sq.=y sq.?

OpenStudy (anonymous):

Is y² there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Yes, that's correct, I believe.

OpenStudy (anonymous):

Then this is your answer. I hope so.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

y sq.

OpenStudy (anonymous):

y

OpenStudy (anonymous):

1/y

OpenStudy (anonymous):

1/y sq.

OpenStudy (anonymous):

-1/y sq.

OpenStudy (anonymous):

Well, I don't know. I've never face this problem before, so I'm not sure exactly how to solve this. Sorry about this.

OpenStudy (anonymous):

they said that the answer was y sq.

OpenStudy (anonymous):

so, do I ignore k sq.?

OpenStudy (anonymous):

if that is the case, 1/x sq.=y sq.

OpenStudy (anonymous):

Yeah, I guess so.

OpenStudy (anonymous):

\[\large y=\frac{1}{x}\] square both sides \[\large y^2 = \frac{1}{x^2}\]

OpenStudy (anonymous):

k